# General Hamiltonian description of nonreciprocal interactions

### v3

@@ C.2.1 Individual-update Glauber scheme: Detailed balance followed by imposing the constraint
$w\left(\theta_{i}\rightarrow\theta_{i}^{\prime}\right)=\frac{1}{2}\left(1-\tanh\frac{H(\theta^{\prime}_{i},\varphi_{i})-H(\theta_{i},\varphi_{i})}{2T}\right)\,.$ (91)
Note the explicit presence of the constraint in Eq. (90), corresponding to Step 2 from the main text. We emphasize that this ad-hoc procedure violates detailed balance – a prerequisite for having a nonequilibrium steady state.
Next, we approximate Eq. (90) for small $\delta$ as
$\Bigg{\{}\left[\rho(\theta_{i},\varphi_{i})+\delta\times\left(\partial_{\theta_{i}}+\partial_{\varphi_{i}}\right)\rho(\theta_{i},\varphi_{i})\right]\frac{1}{2}\left(1-\tanh\frac{-\delta\times\partial_{\theta_{i}}H}{2T}\right)-\rho(\theta_{i},\varphi_{i})\frac{1}{2}\left(1-\tanh\frac{+\delta\times\partial_{\theta_{i}}H}{2T}\right)\Bigg{\}}\Bigg{|}_{\text{constr.}}=0.$
Performing a Taylor expansion on the above equation keeping terms to first order in $\delta$ and multiplying by $T$ yields
$\delta\Big{\{}T\left(\partial_{\theta_{i}}+\partial_{\varphi_{i}}\right)\rho(\theta_{i},\varphi_{i})+\rho(\theta_{i},\varphi_{i})\partial_{\theta_{i}}H\Big{\}}\Big{|}_{\text{constr.}}=0.$ (92)
Since this condition must be satisfied for any $\delta$, we arrive at the following equation for the phase-space distribution function $\rho$
$T\left(\partial_{\theta_{i}}+\partial_{\varphi_{i}}\right)\rho(\theta_{i},\varphi_{i})+\rho(\theta_{i},\varphi_{i})\partial_{\theta_{i}}H\Big{|}_{\text{constr.}}$ $=$ $0$ (93)
Thus, we find that Eq. (93) satisfies Eq. (84) derived above. Therefore, the long-time steady state of the constrained Glauber dynamics reproduces the steady state of the Langevin dynamics.
C.2.2 Joint-update Glauber scheme: Global balance
Let us now discuss an alternative Glauber algorithm, which leads to the same result. Consider again a Markov process where every state $(\theta_{i},\varphi_{i})$ satisfies the constraint $\theta_{i}=\varphi_{i}+\pi$. This time, the constraint is enforced by simultaneously updating the original and auxiliary spins, which change their angles by the same amount $\delta$ as $\theta\rightarrow\theta^{\prime}=\theta+\delta$ and $\varphi\rightarrow\varphi^{\prime}=\varphi+\delta$.
Assuming that the dynamics reaches a nonequilibrium steady state, it satisfies global balance [38]
$\int_{-\pi}^{\pi}\left[\rho(\theta_{i}+\delta,\varphi_{i}+\delta)W(\theta_{i}+\delta,\varphi_{i}+\delta\rightarrow\theta_{i},\varphi_{i})-\rho(\theta_{i},\varphi_{i})W(\theta_{i},\varphi_{i}\rightarrow\theta_{i}+\delta,\varphi_{i}+\delta)\right]\mathrm{d}\delta=0,$ (94)
where the constraint is built-in. Note the additional freedom we have in choosing the precise form of the rates $W$: so long as they satisfy global balance, they can be used to sample the steady-state distribution.
[…]
$w\left(\theta_{i}\rightarrow\theta_{i}^{\prime},\varphi_{i}\rightarrow\varphi_{i}^{\prime}\right)=\frac{1}{2}\left(1-\tanh\left(\frac{H\left(\theta^{\prime},\varphi\right)-H\left(\theta,\varphi^{\prime}\right)}{4T}\right)\right)\,.$ (95)
@@ C.2.2 Joint-update Glauber scheme: Global balance
In other words, we compare the change in energy resulting from the update of each variable $\theta,\varphi$, separately. In particular, we change the angle of the original spin as $\theta\rightarrow\theta^{\prime}{=}\theta+\delta$. The resulting energy difference between the two states is given by $H(\theta^{\prime}_{i},\varphi_{i})-H(\theta_{i},\varphi_{i})$. We also compute the energy difference when the corresponding auxiliary spin is changed under the constraint, i.e., $\varphi\rightarrow\varphi^{\prime}{=}\varphi+\delta$, with the energy difference $H(\theta_{i},\varphi^{\prime}_{i})-H(\theta_{i},\varphi_{i})$. This gives a total change of $H\left(\theta^{\prime},\varphi\right)-H\left(\theta,\varphi^{\prime}\right)$. There is an additional factor of $1/2$ in the argument of the $\tanh$-function, compared to the previous section’s rates, which takes into account the update of both degrees of freedom.
For small $\delta\in[-\Delta,\Delta]$ we can approximate Eq. (94) as
$\int_{-\Delta}^{\Delta}\Bigg{\{}\left[\rho(\theta_{i},\varphi_{i})+\delta\times\left(\partial_{\theta_{i}}+\partial_{\varphi_{i}}\right)\rho(\theta_{i},\varphi_{i})+\frac{1}{2}\delta^{2}\times\left(\partial_{\theta_{i}}+\partial_{\varphi_{i}}\right)^{2}\rho(\theta_{i},\varphi_{i})\right]$
×$\times\frac{1}{2}\left(1-\tanh\frac{-2\left(\partial_{\theta_{i}}H-\partial_{\varphi_{i}}H\right)\times\delta-\left(\partial_{\theta_{i}}^{2}H-\partial_{\varphi_{i}}^{2}H\right)\times\delta^{2}}{8T}\right)$
$-\rho(\theta_{i},\varphi_{i})\frac{1}{2}\left(1+\tanh\frac{-2\left(\partial_{\theta_{i}}H-\partial_{\varphi_{i}}H\right)\times\delta-\left(\partial_{\theta_{i}}^{2}H-\partial_{\varphi_{i}}^{2}H\right)\times\delta^{2}}{8T}\right)\Bigg{\}}\Bigg{|}_{\text{constr.}}\ \mathrm{d}\delta=0.$ (96)
Next, we perform a Taylor expansion on the above equation. The first-order contribution vanishes immediately due to the symmetry of the integrand:
@@
$\frac{1}{2}\Big{\{}T\left(\partial_{\theta_{i}}+\partial_{\varphi_{i}}\right)\rho(\theta_{i},\varphi_{i})+\frac{1}{2}\rho(\theta_{i},\varphi_{i})\left(\partial_{\theta_{i}}H-\partial_{\varphi_{i}}H\right)\Big{\}}\Big{|}_{\text{constr.}}\;\int_{-\Delta}^{\Delta}\delta\;\mathrm{d}\delta=0.$ (97)
@@
$\frac{1}{4}\Big{\{}\left(\partial_{\theta_{i}}+\partial_{\varphi_{i}}\right)\left[T\left(\partial_{\theta_{i}}+\partial_{\varphi_{i}}\right)\rho(\theta_{i},\varphi_{i})+\frac{1}{2}\rho(\theta_{i},\varphi_{i})\left(\partial_{\theta_{i}}H-\partial_{\varphi_{i}}H\right)\right]\Big{\}}\Big{|}_{\text{constr.}}\ \int_{-\Delta}^{\Delta}\delta^{2}\mathrm{d}\delta$ $=$ $0$ (99)
@@
$\left(\partial_{\theta_{i}}+\partial_{\varphi_{i}}\right)\left[T\left(\partial_{\theta_{i}}+\partial_{\varphi_{i}}\right)\rho(\theta_{i},\varphi_{i})+\frac{1}{2}\rho(\theta_{i},\varphi_{i})\left(\partial_{\theta_{i}}H-\partial_{\varphi_{i}}H\right)\right]\Big{|}_{\text{constr.}}\ =0.$ (100)
Multiplying by a test function $f(\theta_{i},\varphi_{i})$ and integrating over phase space, we find that Eq. (100) is equivalent to Eq. (84) since, under the constraint, we have $-\partial_{\varphi_{i}}H=\partial_{\theta_{i}}H$.
Both of these derivations show that the constrained Glauber algorithms proposed here can be used to sample steady states of nonreciprocal systems efficiently. In Figure 6, we test these algorithms against the original Langevin dynamics. The average magnetization and specific heat obtained from Monte-Carlo simulations of constrained Glauber dynamics using the joint update (red) and individual update (blue) exhibit behaviors similar to those obtained from Langevin dynamics. All three methods yield the same critical temperature, $T_{c}/J=0.79\pm 0.03$ (joint update and individual update) and $T_{c}/J=0.78\pm 0.02$ (Langevin) within statistical uncertainty, for the phase transition, cf. Sec. D. Notice that the average magnetization and standard error obtained from these three methods generally agree well, except near the critical temperature, where fluctuations are stronger due to criticality.
Unless explicitly stated otherwise, we use the individual-update Glauber algorithm from Sec. C.2.1 for the numerical simulations in this work.
Figure 6: Average magnetization $\langle m\rangle$ (a) and specific heat $C_{v}$ (b) obtained from Monte-Carlo simulations of constrained Glauber dynamics using the joint update (red, Sec. C.2.2), and individual update (blue, Sec. C.2.1), as well as simulations of the original Langevin dynamics (green) as functions of temperature $T$. All methods yield the same critical temperature within statistical uncertainty, $T_{c}=0.79\pm 0.03J$ (joint update and individual update) and $T_{c}=0.78\pm 0.02J$ (Langevin). The remaining parameter values are the same as in Fig. 3.
Appendix D Determining the critical temperature

### v4

@@ D.2 Glauber dynamics description
$w\left(\theta_{i}\rightarrow\theta_{i}^{\prime},\varphi_{i}\rightarrow\varphi_{i}^{\prime}\right)=\frac{1}{2}\left(1-\tanh\left(\frac{H\left(\theta^{\prime},\varphi\right)-H\left(\theta,\varphi^{\prime}\right)}{4T}\right)\right)\,$, (169)
@@
where the constraint is built-in. In other words, we compare the change in energy resulting from the update of each variable $\theta,\varphi$, separately. In particular, we change the angle of the original spin as $\theta\rightarrow\theta^{\prime}{=}\theta+\delta$. The resulting energy difference between the two states is given by $H(\theta^{\prime}_{i},\varphi_{i})-H(\theta_{i},\varphi_{i})$. Next, we compute the energy difference when the corresponding auxiliary spin is changed under the constraint, i.e., $\varphi\rightarrow\varphi^{\prime}{=}\varphi+\delta$, with the energy difference $H(\theta_{i},\varphi^{\prime}_{i})-H(\theta_{i},\varphi_{i})$. This gives a total change of $H\left(\theta^{\prime},\varphi\right)-H\left(\theta,\varphi^{\prime}\right)$.
Notice the additional factor of $1/2$ in the argument of the $\tanh$-function, which accounts for the update of both degrees of freedom; it serves (i) to derive the correct equation for the phase-space distribution and time-dependent dynamics, cf. Eq. (175) and Eq. (182), respectively, and (ii) to reproduce the correct rates in the reciprocal limit (cf. Sec. D.2.3 below).
The detailed balance condition in Eq. (142) requires that the probability current vanishes on every channel, i.e., $\mathcal{J}(\theta_{i},\varphi_{i}\rightarrow\theta_{i}+\delta,\varphi_{i}+\delta)=0$ in Eq. (168) for every $\delta\in[-\Delta,\Delta]$. However, this condition is generally not satisfied for nonreciprocal systems, which break time-reversal symmetry. Therefore, in Sec D.2.1, we derive the steady-state distribution of the Glauber dynamics based on the global balance and then consider the time-dependent distribution in Sec. D.2.2 when the global balance condition is not satisfied.
D.2.1 Steady-state distribution from global balance
The steady-state distribution is derived from the balance of total probability flux into and out of a given configuration, which constitutes the so-called global balance [47]
@@ D.2.1 Steady-state distribution from global balance
$\sum_{i}\int_{-\pi}^{\pi}\mathrm{d}\delta\left[\mathcal{J}(\theta_{i},\varphi_{i}\rightarrow\theta_{i}+\delta,\varphi_{i}+\delta)-\mathcal{J}(\theta_{i}+\delta,\varphi_{i}+\delta\rightarrow\theta_{i},\varphi_{i})\right]\Big|_{\mathrm{constr.}}=0.$ (170)
In detail, it reads as
$\sum_{i}\int_{-\pi}^{\pi}\mathrm{d}\delta\big[\rho(\theta_{i}+\delta,\{\vec{\theta}\}_{j\neq i},\varphi_{i}+\delta,\{\vec{\varphi}\}_{j\neq i})W(\theta_{i}+\delta,\varphi_{i}+\delta\rightarrow\theta_{i},\varphi_{i})$
$-\rho(\vec{\theta},\vec{\varphi})W(\theta_{i},\varphi_{i}\rightarrow\theta_{i}+\delta,\varphi_{i}+\delta)\big]\Big|_{\text{constr.}}=0.$ (171)
Here, we use $W(\theta_{i}+\delta,\varphi_{i}+\delta\rightarrow\theta_{i},\varphi_{i})=w(\theta_{i}+\delta,\varphi_{i}+\delta\rightarrow\theta_{i},\varphi_{i})\,\mathbf{1}_{[-\Delta,\Delta]}$, where $\mathbf{1}$ is the indicator function which restricts the range of the shift $\delta$ to the interval $\Delta,\Delta]$. In the Glauber scheme, the proposed angle increment $\delta$ is drawn independently at each update step. Consequently, to enforce global balance, one must sum over all lattice sites $i$ and integrate over all proposed shifts $\delta$ in Eq. (D.2.1). This guarantees that we have considered all probability flux channels.
For small $\delta\in[-\Delta,\Delta]$ we can approximate Eq. (D.2.1) as
$\sum_{i}\int_{-\Delta}^{\Delta}\Bigg\{\left[\rho(\theta_{i},\varphi_{i})+\delta\times\left(\partial_{\theta_{i}}+\partial_{\varphi_{i}}\right)\rho(\vec{\theta},\vec{\varphi})+\frac{1}{2}\delta^{2}\times\left(\partial_{\theta_{i}}+\partial_{\varphi_{i}}\right)^{2}\rho(\vec{\theta},\vec{\varphi})\right]$
$\times\frac{1}{2}\left(1-\tanh\frac{-2\left(\partial_{\theta_{i}}H-\partial_{\varphi_{i}}H\right)\times\delta-\left(\partial_{\theta_{i}}^{2}H-\partial_{\varphi_{i}}^{2}H\right)\times\delta^{2}}{8T}\right)$
$-\rho(\vec{\theta},\vec{\varphi})\frac{1}{2}\left(1+\tanh\frac{-2\left(\partial_{\theta_{i}}H-\partial_{\varphi_{i}}H\right)\times\delta-\left(\partial_{\theta_{i}}^{2}H-\partial_{\varphi_{i}}^{2}H\right)\times\delta^{2}}{8T}\right)\Bigg\}\Bigg|_{\text{constr.}}\ \mathrm{d}\delta=0.$ (172)
Next, we perform a Taylor expansion in $\delta$ of the above equation. The first-order contribution vanishes immediately due to the symmetry of the integrand under $\delta\to-\delta$:
$\sum_{i}\frac{1}{2}\Big\{T\left(\partial_{\theta_{i}}+\partial_{\varphi_{i}}\right)\rho(\vec{\theta},\vec{\varphi})+\frac{1}{2}\rho(\vec{\theta},\vec{\varphi})\left(\partial_{\theta_{i}}H-\partial_{\varphi_{i}}H\right)\Big\}\Big|_{\text{constr.}}\;\int_{-\Delta}^{\Delta}\delta\;\mathrm{d}\delta=0.$ (173)
Hence, the steady-state distribution always satisfies global balance to order $\mathcal{O}(\delta)$.
@@
$\sum_{i}\frac{1}{4}\Big\{\left(\partial_{\theta_{i}}+\partial_{\varphi_{i}}\right)\left[T\left(\partial_{\theta_{i}}+\partial_{\varphi_{i}}\right)\rho(\vec{\theta},\vec{\varphi})+\frac{1}{2}\rho(\vec{\theta},\vec{\varphi})\left(\partial_{\theta_{i}}H-\partial_{\varphi_{i}}H\right)\right]\Big\}\Big|_{\text{constr.}}\ \int_{-\Delta}^{\Delta}\delta^{2}\mathrm{d}\delta$ $=$ $0.$ (174)
@@
$\sum_{i}\left(\partial_{\theta_{i}}+\partial_{\varphi_{i}}\right)\left[T\left(\partial_{\theta_{i}}+\partial_{\varphi_{i}}\right)\rho(\vec{\theta},\vec{\varphi})+\rho(\vec{\theta},\vec{\varphi})\partial_{\theta_{i}}H\right]=0.$ (175)
We find that Eq. (175) is equivalent to Eq. (156), after using the relation $\partial_{\varphi_{i}}H=\partial_{\theta_{i}}H$ which is valid under the constraint.
D.2.2 Non-stationary distributions from Monte-Carlo Glauber dynamics
As we have seen, the constrained Glauber dynamics can generate nonequilibrium steady-state distributions through the constraints imposed at each update step. Hence, it is natural to ask whether one can also use Glauber dynamics to obtain the non-stationary states of the original Langevin dynamics. This will enable the application of the Hamiltonian embedding to study non-stationary states of nonreciprocal systems, which can feature configuration-space entropy currents and time-translation symmetry breaking [52]. In particular, we now show that the Glauber dynamics defined by our Hamiltonian embedding reproduces the Fokker–Planck equation derived for the Langevin dynamics in Eq. (161)
We denote the density at iteration step $n$ by $\rho(\vec{\theta},\vec{\varphi},n)$, and recall the corresponding transition rate $w$, defined in Eq. (169). In each update step, the probability of selecting site $j$ is uniformly given by $1/L^{2}$ with $L$ the linear system size; the shift angle $\delta$ is drawn uniformly from the interval $\Delta,\Delta]$, with probability density $1/(2\Delta)$. Sampling under the constraint is reminiscent of directed loop algorithms, which are used to explore the extensive discrete set of ground states in spin models supporting an emergent gauge theory [80], and more efficient algorithms may exist.
After a single Glauber update $n\rightarrow n+1$, the change in the density is given by
$\rho(\vec{\theta},\vec{\varphi},n+1)-\rho(\vec{\theta},\vec{\varphi},n)$ $=\frac{1}{2\Delta L^{2}}\sum_{i}\int_{-\Delta}^{\Delta}\mathrm{d}\delta\Big[\rho(\theta_{i}+\delta,\{\vec{\theta}\}_{i\neq j},\varphi_{i}+\delta,\{\vec{\varphi}\}_{i\neq j},n)\;w(\theta_{i}+\delta\to\theta_{i},\varphi_{i}+\delta\to\varphi_{i})$
$-\rho(\theta_{i},\{\vec{\theta}\}_{i\neq j},\varphi_{i},\{\vec{\varphi}\}_{i\neq j},n)\;w(\theta_{i}\to\theta_{i}+\delta,\varphi_{i}\to\varphi_{i}+\delta)\Big].$ (176)
The right-hand side represents the difference between the total inflow and outflow of density current for the Glauber update. As before, in the limit of small $\Delta$, it can be approximated by expanding both the transition rates and the density up to second order in $\delta$:
$\rho(\vec{\theta},\vec{\varphi},n+1)-\rho(\vec{\theta},\vec{\varphi},n)$ $=\frac{1}{2\Delta L^{2}}\sum_{i}\times$ (177)
@@ D.2.2 Non-stationary distributions from Monte-Carlo Glauber dynamics
$\int_{-\Delta}^{\Delta}\mathrm{d}\delta\Bigg\{\left[\rho(\vec{\theta},\vec{\varphi},n)+\delta\times\left(\partial_{\theta_{i}}+\partial_{\varphi_{i}}\right)\rho(\vec{\theta},\vec{\varphi},n)+\frac{1}{2}\delta^{2}\times\left(\partial_{\theta_{i}}+\partial_{\varphi_{i}}\right)^{2}\rho(\vec{\theta},\vec{\varphi},n)\right]$
$\times\frac{1}{2}\left(1-\tanh\frac{-2\left(\partial_{\theta_{i}}H-\partial_{\varphi_{i}}H\right)\times\delta-\left(\partial_{\theta_{i}}^{2}H-\partial_{\varphi_{i}}^{2}H\right)\times\delta^{2}}{8T}\right)$
$-\rho(\vec{\theta},\vec{\varphi},n)\frac{1}{2}\left(1+\tanh\frac{-2\left(\partial_{\theta_{i}}H-\partial_{\varphi_{i}}H\right)\times\delta-\left(\partial_{\theta_{i}}^{2}H-\partial_{\varphi_{i}}^{2}H\right)\times\delta^{2}}{8T}\right)\Bigg\}\Bigg|_{\text{constr.}}.$
As expected, for stationary distributions where $\rho(\vec{\theta},\vec{\varphi},n)$ is time independent, the equation reduces to Eq. (D.2.1) discussed in Sec. F.4, except for the overall prefactor $1/(2\Delta L^{2})$. Similar to the previous section, we perform a Taylor expansion of the above equation and retain terms up to second order in $\delta$. This yields
$\rho(\vec{\theta},\vec{\varphi},n+1)$ $-\rho(\vec{\theta},\vec{\varphi},n)=\sum_{i}\frac{1}{2\Delta L^{2}}\int_{-\Delta}^{\Delta}\mathrm{d}\delta\Bigg\{\delta\Big[\frac{1}{2}(\partial_{\theta_{i}}+\partial_{\varphi_{i}})\rho(\vec{\theta},\vec{\varphi},n)+\frac{1}{4T}\rho(\vec{\theta},\vec{\varphi},n)(\partial_{\theta_{i}}-\partial_{\varphi_{i}})H\Big]$
$+\delta^{2}\frac{1}{4T}(\partial_{\theta_{i}}+\partial_{\varphi_{i}})\Big[\frac{1}{2}\rho(\vec{\theta},\vec{\varphi},n)(\partial_{\theta_{i}}-\partial_{\varphi_{i}})H+T(\partial_{\theta_{i}}+\partial_{\varphi_{i}})\rho(\vec{\theta},\vec{\varphi},n)\Big]\Bigg\}\Bigg|_{\text{constr.}}.$ (178)
Upon integration, the first-order term in $\delta$ vanishes owing to the symmetry of the integrand. Hence, performing the integration over $\delta$ for the second-order term, we obtain
$\rho(\vec{\theta},\vec{\varphi},n+1)$ $-\rho(\vec{\theta},\vec{\varphi},n)=\sum_{i}\frac{\Delta^{2}}{12L^{2}T}(\partial_{\theta_{i}}+\partial_{\varphi_{i}})\Big[\rho(\vec{\theta},\vec{\varphi},n)\partial_{\theta_{i}}H\big|_{\text{constr.}}+T(\partial_{\theta_{i}}+\partial_{\varphi_{i}})\rho(\vec{\theta},\vec{\varphi},n)\Big]+\mathcal{O}(\Delta^{4}),$ (179)
where we used the equality $\partial_{\theta_{i}}H\big|_{\text{constr.}}=-\partial_{\varphi_{i}}H\big|_{\text{constr.}}$ valid on the constraint manifold. By the same symmetry arguments, the leading order correction is of the order $\mathcal{O}(\Delta^{4})$.
In order to compare the Glauber simulations to real-time Langevin dynamics in practice, we now introduce the notion of Monte-Carlo time, $t_{\mathrm{MC}}$, for the constrained Glauber dynamics. To avoid confusion, the time variable $t$ used outside this section is identified as the Langevin time and is here relabeled as $t_{\mathrm{Lan}}$; this is the proper physical time from the equations of motion of the original nonreciprocal system. We associate one unit of (dimensionless) Monte-Carlo time with updating each spin once (on average), and hence define
$t_{\textrm{MC}}=\frac{n}{L^{2}}.$ (180)
Accordingly, we identify the discrete difference with
$\rho(\vec{\theta},\vec{\varphi},n+1)-\rho(\vec{\theta},\vec{\varphi},n)=\partial_{t_{\mathrm{MC}}}\rho(\theta_{i},\varphi_{i},t_{\mathrm{MC}})\;\mathrm{d}t_{\mathrm{MC}},$ (181)
with $\mathrm{d}t_{\mathrm{MC}}=1/L^{2}$. This allows us to formally define a continuum limit for the Monte-Carlo time.
Following the derivation above, the equation of motion for the density distribution then reads as
$\frac{\partial\rho(\vec{\theta},\vec{\varphi},t_{\textrm{MC}})}{\partial t_{\textrm{MC}}}=\frac{\Delta^{2}}{12T}\sum_{i}(\partial_{\theta_{i}}+\partial_{\varphi_{i}})\Big[\rho\partial_{\theta_{i}}H\big|_{\text{constr.}}+T(\partial_{\theta_{i}}+\partial_{\varphi_{i}})\rho\Big]+\mathcal{O}(\Delta^{4}).$ (182)
This equation is equivalent to the Fokker-Planck equation derived for the Langevin dynamics in Eq. (161), up to the prefactor $\Delta^{2}/(12T)$. In units where the friction coefficient is set to unity, it follows from the equations of motion in Eq. (82) that the interaction strength $J$, and hence energy, has the same unit as frequency (inverse time). Since the only model-independent energy scale in the Glauber rate $w$ is the temperature $T$ (which also has units of energy), we find that the time unit is set by the inverse bath temperature $T$. The dimensionless factor $\Delta^{2}/12$ is a property of the Glauber update rule, and is most likely not universal.
These results demonstrate that the constrained Glauber dynamics reproduce the same non-stationary evolution as the Langevin dynamics as $\Delta\to 0$, provided the Monte-Carlo time $t_{\mathrm{MC}}$ is related to the physical Langevin time $t_{\mathrm{Lan}}$ through a proportionality constant, namely
$\frac{\Delta^{2}}{12T}t_{\mathrm{MC}}=t_{\mathrm{Lan}}.$ (183)
Physically, this relation implies that the lower the bath temperature $T$, the fewer Monte-Carlo updates are necessary to reach the same unit of physical Langevin time at fixed $\Delta$. Moreover, since the constrained Glauber update rule reproduces the nonreciprocal physics in the limit $\Delta\ll 1$ and because the next-order correction in Eq. (182) is $\Delta^{4}$, there is a natural timescale set by $1/\Delta^{2}$ (at a fixed temperature $T$), beyond which Glauber dynamics is expected to disagree with Langevin dynamics; we emphasize that this scale can be controllably extended to an arbitrarily long Langevin time by decreasing $\Delta$ which, however, requires increasing the number of Glauber iterations $n$.
We provide numerical evidence in support of Eq. (183) in Fig. 13 and Sec. G.
D.2.3 Eliminating the auxiliary degree of freedom from the Glauber transition rates
Whereas the nonreciprocal equations of motion of interest depend only on the original degrees of freedom $\theta$, the Glauber update rules derived in the preceding subsections appear to depend on both the original and the auxiliary degrees of freedom: $w\left(\theta_{i}\rightarrow\theta_{i}^{\prime},\varphi_{i}\rightarrow\varphi_{i}^{\prime}\right)$; this is because the Glauber transition rate $w$ was derived using the Hamiltonian embedding. As a result, one might be tempted to argue that the constrained Glauber Monte-Carlo simulations are computationally more expensive than the corresponding original Langevin dynamics.
Here, we show that this is not the case: one can eliminate the auxiliary field $\varphi$ completely from the expression for the constrained Glauber transition rate: $w=w(\theta_{i}\to\theta_{i}^{\prime})$. To get an intuition for why this is the case, notice that, in constrained Glauber dynamics, observables depend only on the original spins $\theta$, while the values of the auxiliary spins $\varphi$ are fixed by the constraint. As a result, the Monte-Carlo fluctuations of $\varphi$ are slaved to those of $\theta$ at each Glauber step.
For clarity, we apply the procedure to the vision-cone XY model using the Hamiltonian in Eq. (141). First, we initialize the simulation in a $(\theta_{i},\varphi_{i})$-configuration that obeys the constraint. Hence, at the start of each Glauber update step, the constraint $\varphi_{i}=\theta_{i}+\pi$ holds, so the auxiliary angle can be expressed in terms of $\theta_{i}$. We then update $\theta_{i}\to\theta_{i}^{\prime}=\theta_{i}+\delta$ and $\varphi_{i}\to\varphi_{i}^{\prime}=\varphi_{i}+\delta$, which preserves the constraint manifold: $\varphi_{i}^{\prime}=\theta_{i}^{\prime}+\pi$. The key observation is that $\varphi_{i}^{\prime}$ may be replaced by $\theta_{i}^{\prime}+\pi$ when evaluating the energy differences which define the Glauber transition rate.
To see this in practice, we will revisit the expression for the transition rate from Eq. (169). The energy difference in the argument of the hyperbolic tangent can be expressed as
$\Delta E$ $=$ $\frac{1}{2}\big(H(\theta^{\prime},\varphi)-H(\theta,\varphi^{\prime})\big)\big|_{\varphi^{\prime}=\theta^{\prime}+\pi}^{\varphi=\theta+\pi}$ (184)
$=$ $\frac{1}{2}\big(H(\theta^{\prime},\varphi)-H(\theta,\varphi)+H(\theta,\varphi)-H(\theta,\varphi^{\prime})\big)\big|_{\varphi^{\prime}=\theta^{\prime}+\pi}^{\varphi=\theta+\pi}$
$=$ $\frac{1}{2}\left(\Delta E_{1}\big|_{\varphi=\theta+\pi}-\Delta E_{2}\big|_{\varphi^{\prime}=\theta^{\prime}+\pi}^{\varphi=\theta+\pi}\right),$
where, crucially, the constraint is applied only after the differences are evaluated.
Now consider the term $\Delta E_{1}$, and note that the energy change associated with updating the original spin $\theta_{i}$, while keeping $\varphi_{i}$ fixed, is given by
$\Delta E_{1}$ $=H(\theta^{\prime},\varphi)-H(\theta,\varphi)$
$=-\sum_{j\in\langle ij\rangle}\Big\{[J_{ij}(\theta_{i}^{\prime})+J_{ji}(\theta_{j})]\cos(\theta_{i}^{\prime}-\theta_{j})-[J_{ij}(\theta_{i})+J_{ji}(\theta_{j})]\cos(\theta_{i}-\theta_{j})\Big\}$
$\quad-\sum_{j\in\langle ij\rangle}\Big\{[J_{ij}(\theta_{i}^{\prime})-J_{ij}(\theta_{i})]\cos(\theta_{j}-\varphi_{i})+J_{ji}(\theta_{j})[\cos(\theta_{i}^{\prime}-\varphi_{j})-\cos(\theta_{i}-\varphi_{j})]\Big\}.$ (185)
Since the Glauber update depends only on the energy difference, the auxiliary angles can be eliminated using the constraint $\varphi=\theta+\pi$, without changing the transition rate. Substituting $\varphi=\theta+\pi$ gives
$\Delta E_{1}\big|_{\varphi=\theta+\pi}$ $=H(\theta^{\prime},\theta+\pi)-H(\theta,\theta+\pi)$
$=-\sum_{j\in\langle ij\rangle}\Big\{[J_{ij}(\theta_{i}^{\prime})+J_{ji}(\theta_{j})]\cos(\theta_{i}^{\prime}-\theta_{j})-[J_{ij}(\theta_{i})+J_{ji}(\theta_{j})]\cos(\theta_{i}-\theta_{j})\Big\}$
$\quad+\sum_{j\in\langle ij\rangle}[J_{ij}(\theta_{i}^{\prime})-J_{ij}(\theta_{i})]\cos(\theta_{j}-\theta_{i})+\sum_{j\in\langle ij\rangle}J_{ji}(\theta_{j})[\cos(\theta_{i}^{\prime}-\theta_{j})-\cos(\theta_{i}-\theta_{j})]$
$=-\sum_{j\in\langle ij\rangle}J_{ij}(\theta_{i}^{\prime})[\cos(\theta_{i}^{\prime}-\theta_{j})-\cos(\theta_{i}-\theta_{j})].$ (186)
Likewise, the energy difference associated with updating the auxiliary spin $\varphi_{i}$ while keeping $\theta_{i}$ fixed reads as
$\Delta E_{2}=H(\theta,\varphi^{\prime})-H(\theta,\varphi)=-\sum_{j\in\langle ij\rangle}J_{ij}(\theta_{i})[\cos(\theta_{j}-\varphi_{i}^{\prime})-\cos(\theta_{j}-\varphi_{i})].$ (187)
Applying the constraint by substituting $\varphi^{\prime}=\theta^{\prime}+\pi$ and $\varphi=\theta+\pi$ yields
$\Delta E_{2}\big|_{\varphi^{\prime}=\theta^{\prime}+\pi}^{\varphi=\theta+\pi}=H(\theta,\theta^{\prime}+\pi)-H(\theta,\theta+\pi)=\sum_{j\in\langle ij\rangle}J_{ij}(\theta_{i})[\cos(\theta_{i}^{\prime}-\theta_{j})-\cos(\theta_{i}-\theta_{j})].$ (188)
Therefore, the total energy difference entering the transition rate in Eq. (169) simplifies to
$\Delta E$ $=\frac{1}{2}\left(\Delta E_{1}\big|_{\varphi=\theta+\pi}-\Delta E_{2}\big|_{\varphi^{\prime}=\theta^{\prime}+\pi}^{\varphi=\theta+\pi}\right)$
$=-\frac{1}{2}\sum_{j\in\langle ij\rangle}[J_{ij}(\theta_{i}^{\prime})+J_{ij}(\theta_{i})][\cos(\theta_{i}^{\prime}-\theta_{j})-\cos(\theta_{i}-\theta_{j})].$ (189)
Consequently, the joint-update constrained Glauber dynamics can be implemented using only the original variables $\theta_{i}$, and the corresponding transition rate is
$w(\theta_{i}\to\theta_{i}^{\prime})=\frac{1}{2}\left(1-\tanh\frac{\Delta E}{2T}\right),$ (190)
with $\Delta E$ given by Eq. (189).
Notably, in the reciprocal limit $\Psi{=}\pi$, the interaction becomes orientation-independent: $J_{ij}(\theta_{i}^{\prime})=J_{ij}(\theta_{i})=J_{0}$. Then $\Delta E=-J_{0}\sum_{j\in\langle ij\rangle}\big[\cos(\theta_{i}^{\prime}-\theta_{j})-\cos(\theta_{i}-\theta_{j})\big]$, which matches exactly the energy difference obtained from the conventional XY Hamiltonian $H_{\mathrm{XY}}=-J_{0}\sum_{\langle ij\rangle}\cos(\theta_{i}-\theta_{j})$ (and hence justifies the factor of $1/2$ in the definition of $\Delta E$ above a posteriori). Thus, although imposing the constraint breaks detailed balance in general, the latter is restored in the reciprocal limit, as expected.
As a consequence of eliminating the auxiliary degrees of freedom $\varphi$, the total number of XY-spins to be simulated in the constrained Glauber algorithm is equivalent to that used in the original Langevin dynamics. Note that we proved this elimination procedure only for the constrained Glauber dynamics considered in this work; further research is required to determine whether, and under which conditions, it extends to other energy-difference-based algorithms.
Finally, let us emphasize that the role of the Hamiltonian embedding here is to provide the correct Glauber transition rate (which we have proved to give rise to the correct Fokker-Planck equation in the limit of $\delta\to 0$, cf. Sec. F.4).
Appendix E Determining the critical temperature for the vision-cone model